# Did I Vet You Before? Assessing the Chrome Web Store Vetting Process through Browser Extension Similarity

### v1

@@ 4.1. Feature Extraction
Manifest. We flatten the properties found in the extension manifest to convert them from a nested JSON object to key-value pairs. We exclude the author, name, short_name, description, version, key and update_url properties as they typically do not provide reliable information about the behavior of an extension and, in the worst case, could help the attacker to trick the model into giving a different embedding to the extension. We also identify and enumerate all code entrypoints found in the manifest. Per Google’s documentation (Google Chrome Developers, 2012a), we consider any of the following items to be valid entrypoints: $(i)$ content scripts; $(ii)$ background pages or service workers; $(iii)$ popup pages, i.e., browser actions; and $(iv)$ overridden, DevTools, side panel, or options pages.
@@
Static Analysis. We statically analyze the JavaScript code of entrypoints. We use esbuild (Wallace and Open-source contributors, 2024) for module resolution, bundling, tree shaking (i.e., dead code removal) and minification. We then use Babel (Ng et al., 2024), a well-maintained and widely used tool in the modern web ecosystem, to generate the Abstract Syntax Tree (AST) of the resulting bundle. Although previous work has relied on Esprima (Hidayat and Open-source contributors, 2021) for this purpose (Fass et al., 2021; Pantelaios et al., 2020; Solomos et al., 2022; Somé, 2019; Eriksson et al., 2022), we decided to use a more modern alternative given that Esprima has not received any major updates since 2018. Lastly, we traverse the AST to extract the list of API calls as relevant features to model the behavior of the extension. We define an API call as an invocation to an Extension API (Google Chrome Developers, 2024a) from the chrome or browser object, or to a Web API (MDN contributors, 2024b) within the navigator object.
@@
Dynamic Analysis. To overcome the limitations of static analysis and detect API calls we might otherwise miss (e.g., code being executed in a call to eval), we complement our feature extraction pipeline with a dynamic analysis stage. We use isolated-vm (Laverdet and Open-source contributors, 2024) to create a sandboxed environment that supports ES6 modules to safely run untrusted JavaScript code. As this environment lacks the webpage DOM and Web and Browser APIs, we mock any objects accessed by the extension at runtime to prevent it from crashing for as long as possible. We monitor calls to mocked objects and log API calls as in static analysis. While not as accurate as running the extension in an actual web browser, this approach is more scalable, as it takes a few seconds instead of minutes to analyze an extension. In addition, this environment relies on the same JavaScript engine that Chromium uses (The Chromium Authors, 2024), guaranteeing accurate output for code snippets that depend only on the ECMAScript specification (i.e., without invoking APIs that will be mocked by our sandbox).
4.2. Vectorization
@@
Inspired by Zero-Shot Learning (ZSL) (Pushp and Srivastava, 2017), we use Sentence-Transformers (Reimers and Gurevych, 2019) with the “all-distilroberta-v1” variant of the DistilRoBERTa model111 Available at https://huggingface.co/sentence-transformers/all-distilroberta-v1 to compute 768-dimensional embeddings of the generated documents. DistilRoBERTa is a general purpose model based on the BERT framework that was trained on a large and diverse corpus of English texts (Liu et al., 2019). It uses knowledge distillation to improve performance while retaining a high accuracy (Sanh et al., 2019). We found this model to be a good fit for our needs as it achieves a good balance between sentence performance and encoding speed.222 See comparison at https://www.sbert.net/docs/pretrained_models.html Using the default tokenizer provided with the model, the average length of our documents is 300 tokens, with a standard deviation of 228 tokens. While some inputs are truncated after exceeding the maximum sequence length of 512 tokens set by the model, this number is fairly small as 86% of the documents fit within the limit.
4.3. Clustering
We employ density-based clustering to group together extensions with a very similar behavior based on their embeddings, and to filter out outliers that do not belong to any cluster. Prior to the clustering, we use Principal Component Analysis (PCA) to reduce the dimensionality of the standardized embeddings to 161 components for performance, keeping 95% of the amount of variance. Since DistilRoBERTa already outputs normalized embeddings, we skip the standardization step. Then, we use Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) (McInnes et al., 2017). We choose HDBSCAN over other clustering algorithms like K-Means as it does not require knowing the number of clusters beforehand. We pick a minimum cluster size of 5 data points and lower the minimum samples parameter to 2. This configuration should discard points with low stability (i.e., outliers) while still allowing us to create clusters in less dense areas.
Visualization. To facilitate the visualization of the extensions in our dataset and their similarities, we use Uniform Manifold Approximation and Projection (UMAP) (McInnes et al., 2018) to project the high-dimensional space formed by the standardized extension embeddings onto a 2D plane. This step does not affect the results of the clustering process, as we rely solely on HDBSCAN for assigning cluster IDs to extensions.
4.4. Evaluation
@@
Ground Truth. Given the lack of previous work in this area, we could not find a dataset of similar CRX files, nor a tool to facilitate this process. Therefore, we create our own ground truth by listing popular extensions in the CWS and using the embeddings generated by our pipeline to find similar versions of these items. Additionally, we take advantage of developers publishing the same extension across stores to add them to the list of similar pairs. That is, we manually look for samples of featured CWS extensions that are also published on Edge Add-ons (Microsoft Edge Developers, 2023) or Opera add-ons (Opera Developers, 2023), as we expect those to be identical. We build pairs of different extensions by looking for items that offer the same functionality but implement it differently. For example, uBlock Origin and Privacy Badger make a pair of different ad blockers.
@@
For a fairer evaluation, we complement the ground truth with manually-reviewed pairs of similar and different extensions taken from random clusters obtained after building the pipeline. We provide the resulting list of pairs as an artifact for reproducibility (see Section 1).
Results. Using the process described above, we managed to put together a ground truth dataset of 110 pairs of similar extension versions, and another 110 pairs of different ones. When validating against the comprised ground truth, SimExt shows an accuracy of 83.2%, with a precision of 85.4% and a recall of 80.0%. We consider these results adequate for the task: as the first work to use embeddings for clustering browser extensions, the high accuracy obtained gives us confidence in the proposed novel methodology and positions us well to conduct the first examination of the CWS vetting process.
5. Effectiveness of the Vetting Process
@@
5.1. Infringing Extensions Detection
As discussed in Section 4, we assume that extensions that are very similar to vetted extensions are most likely in violation of store policies and should likely be taken down. We acknowledge that this is a strong assumption, and we validate it in Section 6 by performing a manual analysis of the clusters and providing compelling examples of nearly identical extensions found by SimExt. To benchmark how good the CWS is at finding these potentially infringing extensions, we quantify how many extensions that are still published as of the latest snapshot of our dataset belong to a vetted cluster. We define an infringing cluster as a set of 2 or more similarly behaving extensions in which at least one of them has been taken down by Google (i.e., has been vetted) according to the the labels they provide, as presented in Section 3.
@@ 5.1. Infringing Extensions Detection
Figure 4. Scatter plots of extensions belonging to an infringing cluster. Taken down (vetted) extensions in red, unpublished by the developer in green, and still published in blue.
Figure 4 shows two-dimensional projections of the embeddings of extensions found in infringing clusters. In total, our tool finds 6,444 infringing clusters comprised of 100,957 extensions. These clusters contain 72,699 extensions that were taken down (72%), and an additional smaller—yet significant—set of 11,237 unpublished extensions (11%). According to CWS labels, the aforementioned were voluntarily unpublished by their developers without Google raising a policy violation despite being similar to vetted extensions. In Section 8 we discuss why we believe publishers might prefer to intentionally remove their extensions over receiving a takedown by Google. Lastly, our pipeline finds 17,021 published extensions that account for the remaining 17% of infringing clusters. Overall, these numbers are reasonably good and show that the CWS has some capability to detect infringing extensions. Nevertheless, we reiterate that the number of published infringing items is an estimate derived from our pipeline and the actual amount may vary.
@@
Repeat Offenders. We find evidence of 6,156 repeat offenders, or publishers with multiple vetted extensions. Up to 11% of these publishers (663 accounts) have at least one published infringing extension as of our last date of crawling. Put another way, 11% of the published infringing extensions (1,896 items) come from repeat offenders. Another takeaway is that these accounts were not banned from the store after repeatedly uploading confirmed infringing extensions. According to the CWS, “repeated violations” of the Program Policies will result in the suspension of the developer account (Google Chrome Developers, 2022d). The existence of known repeat offenders with published extensions is a strong indication that this policy is not being properly enforced.
@@
5.2. Survival Analysis
We perform a survival analysis (Clark et al., 2003) on infringing extensions to assess how rapidly the CWS removes policy-violating content. To determine how long these extensions remain published in the store, we measure the lifetime or duration in days between the release date of the infringing extension version and its “death” (removal date). For published infringing extensions—for which we do not observe this event—, we take the latest date of our crawling as the removal date, and label them as right-censored observations.
@@ 5.2. Survival Analysis
Figure 5. Kaplan–Meier curves of the survival of infringing extensions by vetting label.
To study the likelihood of survival for all infringing items, we use the lifelines Python library (Davidson-Pilon, 2019) to plot Kaplan–Meier curves and assist with related statistical analysis. When plotting these curves using all 100k infringing extensions, the confidence intervals became too narrow and were not visible in the figure. Since this makes it difficult to assess the uncertainty around the survival estimates, we perform several experiments with different random sample sizes to find the best visualization. Seeing that the curves have no meaningful visible changes other than the confidence intervals, we take a random sample of 1,000 extensions (per variable or class) to display the Kaplan–Meier curves in a clearer manner. The curve labeled as “All” from Figure 5 shows the survival probability of infringing extensions. The median survival time is 581 days, confirming our earlier assertion that more than half of vetted extensions remain in the CWS for more than a year. However, while the distribution of vetted extensions shows few items with a life expectancy greater than a decade, the Kaplan–Meier estimate predicts that 6% of infringing extensions will surpass the same period. This result makes sense due to the inclusion of published infringing extensions as censored observations in the calculation of this curve.
@@
6. Characterizing Infringing Extensions
Table 1. Top 65 largest infringing clusters. Includes only the first 15 NTE clusters, skipping the rest to provide a more representative sample. Footer row has totals from all clusters, including skipped. Highlighted rows contain published extensions.
@@ 60 Boilerplate aemnapldoeccnbmfmifibgdjimkdgnig League of Legends Acc… 210 76 16 108 0 160 28k 4y 5y
T: Total   P: Published   PV: Policy Violation   MPV: Minor Policy Violation   M: Malware
This section examines what kinds of infringing extensions are the most prevalent in our dataset (RQ2). For this analysis, we take the infringing clusters obtained in Section 5 and sort them by descending size, prioritizing those with the most extensions. That is, we are interesting in finding the most common groups of behavior across infringing extensions. We select the top largest infringing clusters and assign a kind or type of content to each one by manually inspecting a random sample of the extensions it contains.
Table 1 shows the ranked list of the top 65 infringing clusters. Almost all of these clusters contain NTEs. To provide a broader picture of what kind of extensions lie in these clusters, we only display up to 15 NTE clusters in Table 1, allowing for other kinds of clusters to appear. Therefore, rows that are missing from the table, such as cluster #18, are omitted because they also contain NTEs.
Top infringing clusters range in size from hundreds to a few thousand extensions. The number of unique publishers per cluster is always lower than its size, confirming that some developers republish the same extension with the same account. Out of the top 65 infringing clusters, containing 44k extensions, we only find 6.2k publishers. With some permissible exceptions allowed by the CWS, this is a strong indication of repetitive content (Google Chrome Developers, 2020c). NTE clusters have the highest removal rate, with almost no extensions still published at the end of our crawl. Interestingly, we find some instances of no consensus over the vetting label among extensions in the same cluster. For example, clusters #2 and #4 contain almost identical extensions that have been removed completely from the CWS, yet the reason varies among Minor Policy Violation, Policy Violation, and Malware. There is a high variance in the cluster impact, finding some with millions of affected users. The most severe case is cluster #12, which lasted just 2 months in the store but managed to gather a total of over 150 million users. In terms of lifetime, top infringing extensions last an average of one year. Some clusters like #34 are even fairly new, having been created a year ago and containing infringing extensions that have not yet been removed from the CWS.
@@ 6.1. New Tab Extensions (NTEs)
To prevent publishers from flooding the CWS with low-quality, repetitive extensions, Google released new program policies in June 2021 that prohibited “multiple extensions with highly similar functionality,” specifically citing “wallpaper extensions” as an example (Soares and Ackerman, 2021). Consequently, close to all NTEs found in the top infringing clusters have already been removed. We hypothesize that this high removal rate, in contrast to non-NTE clusters, is due to the fact that NTEs are easily findable because they include recurring keywords such as “wallpaper” or “new tab” in their names, facilitating the vetting process. Nevertheless, we still find 55 published NTEs in top infringing clusters, and at least 571 in the entire store (i.e., including clusters outside this ranking).
(a) Mountain Wallpaper (coolthemestores.com, 2024)
(b) Salads Wallpaper (wallpaperaddons.com, 2019)
Figure 6. Sample screenshots of published NTEs.
@@
6.2. No Content Extensions
For a extension to be considered eligible for submission to the CWS, it must have a name, description, version, and an icon declared in its manifest (Google Chrome Developers, 2023). Despite satisfying all the minimum requirements, an extension with only these attributes would be useless due to its lack of functionality. Furthermore, according to the Program Policies, it would violate the Minimum Functionality clause by providing no utility whatsoever to the CWS catalog (Google Chrome Developers, 2022c).

### v2

@@ IV-A Feature Extraction
Manifest. We flatten the properties found in the extension manifest to convert them from a nested JSON object to key-value pairs. We exclude the author, name, short_name, description, version, key and update_url properties as they typically do not provide reliable information about the behavior of an extension and, in the worst case, could help the attacker to trick the model into giving a different embedding to the extension. We also identify and enumerate all code entrypoints found in the manifest. Per Google’s documentation [10], we consider any of the following items to be valid entrypoints: $(i)$ content scripts; $(ii)$ background pages or service workers; $(iii)$ popup pages, i.e., browser actions; and $(iv)$ overridden, DevTools, side panel, or options pages.
@@
Static Analysis. We statically analyze the JavaScript code of entrypoints. We use esbuild [63] for module resolution, bundling, tree shaking (i.e., dead code removal) and minification. We then use Babel [43], a well-maintained and widely used tool in the modern web ecosystem, to generate the Abstract Syntax Tree (AST) of the resulting bundle. Although previous work has relied on Esprima [25] for this purpose [9, 48, 58, 59, 8], we decided to use a more modern alternative given that Esprima has not received any major updates since 2018. Lastly, we traverse the AST to extract the list of API calls as relevant features to model the behavior of the extension. We define an API call as an invocation to an Extension API [21] from the chrome or browser object, or to a Web API [40] within the navigator object. In particular, we use the Babel Traverse module and the ReferencedIdentifier and BinaryExpression nodes to find the names of global variables (e.g., with no bindings except for the global scope) being invoked in a program.
@@
Dynamic Analysis. To overcome the limitations of static analysis and detect API calls we might otherwise miss (e.g., code being executed in a call to eval), we complement our feature extraction pipeline with a dynamic analysis stage. We develop Fakeium, our own sandbox environment based on isolated-vm [32] to safely run untrusted JavaScript code. Fakeium intentionally lacks the webpage DOM and Web and Browser APIs. Instead, it works by mocking any objects accessed by the extension at runtime to prevent it from crashing for as long as possible. This mocking is performed by hijacking the global scope using a Proxy object [38]. We monitor calls to mocked objects and log API calls as in static analysis. To ensure that we cover as many code paths as possible, we follow an approach inspired by concolic execution, where we invoke all callbacks and functions found during the analysis, whilst still respecting the original conditional predicates. This approach enables Fakeium to run without the need for an automation tool or monkey to trigger behaviors. While not as accurate as running the extension in an actual web browser, this approach is more scalable, taking just a few seconds instead of minutes to analyze an extension. In addition, this environment relies on the same JavaScript engine that Chromium uses [60], guaranteeing accurate output for code snippets that depend only on the ECMAScript specification (i.e., without invoking APIs that will be mocked by our sandbox). Fakeium is released under the MIT license and distributed as a Node.js package [42].
IV-B Vectorization
@@
Inspired by Zero-Shot Learning (ZSL) [52], we use Sentence-Transformers [54] with the “all-distilroberta-v1” variant of the DistilRoBERTa model222 See https://huggingface.co/sentence-transformers/all-distilroberta-v1 to compute 768-dimensional embeddings of the generated documents. DistilRoBERTa is a general purpose model based on the BERT framework that was trained on a large and diverse corpus of English texts [34]. It uses knowledge distillation to improve performance while retaining a high accuracy [56]. We found this model to be a good fit for our needs as it achieves a good balance between sentence performance and encoding speed.333 See comparison at https://www.sbert.net/docs/pretrained_models.html Using the default tokenizer provided with the model, the average length of our documents is 300 tokens, with a standard deviation of 228 tokens. While some inputs are truncated after exceeding the maximum sequence length of 512 tokens set by the model, this number is fairly small as 86% of the documents fit within the limit.
IV-C Clustering
We employ density-based clustering to group together extensions with a very similar behavior based on their embeddings, and to filter out outliers that do not belong to any cluster. Prior to the clustering, we use Principal Component Analysis (PCA) to reduce the dimensionality of the standardized embeddings for performance while retaining 95% of the amount of variance. This results in PCA deciding to keep 161 components. Since DistilRoBERTa already outputs normalized embeddings, we skip the standardization step. Then, we use Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) [36]. We choose HDBSCAN over other clustering algorithms like K-Means as it does not require knowing the number of clusters beforehand. We pick a minimum cluster size of 5 data points and lower the minimum samples parameter to 2. This configuration should discard points with low stability (i.e., outliers) while still allowing us to create clusters in less dense areas.
Visualization. To facilitate the visualization of the extensions in our dataset and their similarities, we use Uniform Manifold Approximation and Projection (UMAP) [37] to project the high-dimensional space formed by the standardized extension embeddings onto a 2D plane. This step does not affect the results of the clustering process, as we rely solely on HDBSCAN for assigning cluster IDs to extensions.
IV-D Evaluation
@@
Ground Truth. Given the lack of previous work in this area, we could not find a dataset of similar CRX files, nor a tool to facilitate this process. Therefore, we create our own ground truth by listing popular extensions in the CWS and using the embeddings generated by our pipeline to find similar versions of these items. Additionally, we take advantage of developers publishing the same extension across stores to add them to the list of similar pairs. That is, we manually look for samples of featured CWS extensions that are also published on Edge Add-ons [41] or Opera add-ons [45], as we expect those to be identical. We build pairs of different extensions by looking for items that offer the same functionality but implement it differently. For example, uBlock Origin and Privacy Badger make a pair of different ad blockers.
@@
For a fairer evaluation, we complement the ground truth with manually-reviewed pairs of similar and different extensions taken from random clusters obtained after building the pipeline. We provide the resulting list of pairs as an artifact for reproducibility (see Section I).
Results. Using the process described above, we managed to put together a ground truth dataset of 110 pairs of similar extension versions, and another 110 pairs of different ones (220 unique pairs in total). When validating against the comprised ground truth, SimExt shows an accuracy of 83.2%, with a precision of 85.4% and a recall of 80.0%. We consider these results adequate for the task: as the first work to use embeddings for clustering browser extensions, the high accuracy obtained gives us confidence in the proposed novel methodology and positions us well to conduct the first examination of the CWS vetting process.
V Effectiveness of the Vetting Process
@@
V-A Infringing Extensions Detection
As discussed in Section IV, we assume that extensions that are very similar to vetted extensions are most likely in violation of store policies and should likely be taken down. We acknowledge that this is a strong assumption, and we validate it in Section VI by performing a manual analysis of the clusters and providing compelling examples of nearly identical extensions found by SimExt. To benchmark how good the CWS is at finding these potentially infringing extensions, we quantify how many extensions that are still published as of the latest snapshot of our dataset belong to a vetted cluster. We define an infringing cluster as a set of 2 or more similarly behaving extensions in which at least one of them has been taken down by Google (i.e., has been vetted) according to the the labels they provide, as presented in Section III.
@@ V-A Infringing Extensions Detection
Figure 4: Scatter plots of extensions belonging to an infringing cluster. Taken down (vetted) extensions in red, unpublished by the developer in green, and still published in blue.
Figure 4 shows two-dimensional projections of the embeddings of extensions found in infringing clusters. In total, our tool finds 6,444 infringing clusters comprised of 100,957 extensions. These clusters contain 72,699 extensions that were taken down (72%), and an additional smaller—yet significant—set of 11,237 unpublished extensions (11%). According to CWS labels, the aforementioned were voluntarily unpublished by their developers without Google raising a policy violation despite being similar to vetted extensions. In Section VIII we discuss why we believe publishers might prefer to intentionally remove their extensions over receiving a takedown by Google. Lastly, our pipeline finds 17,021 published extensions that account for the remaining 17% of infringing clusters. Overall, these numbers are reasonably good and show that the CWS has some capability to detect infringing extensions. Nevertheless, we reiterate that the number of published infringing items is an estimate derived from our pipeline and the actual amount may vary.
@@
Repeat Offenders. We find evidence of 6,156 repeat offenders, or publishers with multiple vetted extensions. Up to 11% of these publishers (663 accounts) have at least one published infringing extension as of our last date of crawling. Put another way, 11% of the published infringing extensions (1,896 items) come from repeat offenders. Another takeaway is that these accounts were not banned from the store after repeatedly uploading confirmed infringing extensions. According to the CWS, “repeated violations” of the Program Policies will result in the suspension of the developer account [19]. The existence of known repeat offenders with published extensions is a strong indication that this policy is not being properly enforced.
@@
V-B Survival Analysis
We perform a survival analysis [4] on infringing extensions to assess how rapidly the CWS removes policy-violating content. To determine how long these extensions remain published in the store, we measure the lifetime or duration in days between the release date of the infringing extension version and its “death” (removal date). For published infringing extensions—for which we do not observe this event—, we take the latest date of our crawling as the removal date, and label them as right-censored observations.
@@ V-B Survival Analysis
Figure 5: Kaplan–Meier curves of the survival of infringing extensions by vetting label.
To study the likelihood of survival for all infringing items, we use the lifelines Python library [6] to plot Kaplan–Meier curves and assist with related statistical analysis. When plotting these curves using all 100k infringing extensions, the confidence intervals became too narrow and were not visible in the figure. Since this makes it difficult to assess the uncertainty around the survival estimates, we perform several experiments with different random sample sizes to find the best visualization. Seeing that the curves have no meaningful visible changes other than the confidence intervals, we take a random sample of 1,000 extensions (per variable or class) to display the Kaplan–Meier curves in a clearer manner. The curve labeled as “All” from Figure 5 shows the survival probability of infringing extensions. The median survival time is 581 days, confirming our earlier assertion that more than half of vetted extensions remain in the CWS for more than a year. However, while the distribution of vetted extensions shows few items with a life expectancy greater than a decade, the Kaplan–Meier estimate predicts that 6% of infringing extensions will surpass the same period. This result makes sense due to the inclusion of published infringing extensions as censored observations in the calculation of this curve.
@@
VI Characterizing Infringing Extensions
TABLE I: Top 65 largest infringing clusters. Includes only the first 15 NTE clusters, skipping the rest to provide a more representative sample. Footer row has totals from all clusters, including skipped. Highlighted rows contain published extensions. Extension labels are Total (T), Published (P), Policy Violation (PV), Minor Policy Violation (MPV), and Malware (M).
@@ 60 Boilerplate aemnapldoeccnbmfmifibgdjimkdgnig League of Legends Acc… 210 76 16 108 0 160 28k 4y 5y
This section examines what kinds of infringing extensions are the most prevalent in our dataset (RQ2). For this analysis, we take the infringing clusters obtained in Section V and sort them by descending size, prioritizing those with the most extensions. That is, we are interesting in finding the most common groups of behavior across infringing extensions. We select the top largest infringing clusters and assign a kind or type of content to each one by manually inspecting a random sample of the extensions it contains.
Table I shows the ranked list of the top 65 infringing clusters. Almost all of these clusters contain NTEs. To provide a broader picture of what kind of extensions lie in these clusters, we only display up to 15 NTE clusters in Table I, allowing for other kinds of clusters to appear. Therefore, rows that are missing from the table, such as cluster #18, are omitted because they also contain NTEs.
Top infringing clusters range in size from hundreds to a few thousand extensions. The number of unique publishers per cluster is always lower than its size, confirming that some developers republish the same extension with the same account. Out of the top 65 infringing clusters, containing 44k extensions, we only find 6.2k publishers. With some permissible exceptions allowed by the CWS, this is a strong indication of repetitive content [15]. NTE clusters have the highest removal rate, with almost no extensions still published at the end of our crawl. Interestingly, we find some instances of no consensus over the vetting label among extensions in the same cluster. For example, clusters #2 and #4 contain almost identical extensions that have been removed completely from the CWS, yet the reason varies among Minor Policy Violation, Policy Violation, and Malware. There is a high variance in the cluster impact, finding some with millions of affected users. The most severe case is cluster #12, which lasted just 2 months in the store but managed to gather a total of over 150 million users. In terms of lifetime, top infringing extensions last an average of one year. Some clusters like #34 are even fairly new, having been created a year ago and containing infringing extensions that have not yet been removed from the CWS.
@@ VI-A New Tab Extensions (NTEs)
To prevent publishers from flooding the CWS with low-quality, repetitive extensions, Google released new program policies in June 2021 that prohibited “multiple extensions with highly similar functionality,” specifically citing “wallpaper extensions” as an example [57]. Consequently, close to all NTEs found in the top infringing clusters have already been removed. We hypothesize that this high removal rate, in contrast to non-NTE clusters, is due to the fact that NTEs are easily findable because they include recurring keywords such as “wallpaper” or “new tab” in their names, facilitating the vetting process. Nevertheless, we still find 55 published NTEs in top infringing clusters, and at least 571 in the entire store (i.e., including clusters outside this ranking).
(a) Mountain Wallpaper [5]
(b) Salads Wallpaper [64]
Figure 6: Sample screenshots of published NTEs.
@@
VI-B No Content Extensions
For a extension to be considered eligible for submission to the CWS, it must have a name, description, version, and an icon declared in its manifest [20]. Despite satisfying all the minimum requirements, an extension with only these attributes would be useless due to its lack of functionality. Furthermore, according to the Program Policies, it would violate the Minimum Functionality clause by providing no utility whatsoever to the CWS catalog [18].